# ZTF SN Ia DR2: Overview

### v1

@@ 4 Overview of the released data
Details concerning the DR2 data and the derived parameters are given in Smith et al.. This section summarizes the top-level information.
@@ 4.1 Light curves
In this release, we provide , and ZTF light curves for all our 3628 SNe Ia acquired by the ZTF camera installed on the P48. Additional photometric data points, as these obtained from the SEDmachine in camera-mode, are not included in this release.
@@
On the 2960 SNe Ia passing the good sampling criteria, only 9 do not have band and only 2 do not have band detections in the rest-frame phase range of 10 to +40 with respect to maximum light. However, only 46% of our targets have -band detections in this same phase range. Most targets acquired since early 2021 shall have coverage in all three bands as we updated our observing strategy then but are not part of this sample. Future releases of DR2 targets will include additional band observations, after the remaining reference images are processed. For these 2960 targets, we typically have 40 detections (median) in the days phase range, with medians of 9 and 27 points pre- and post-maximum light, respectively. This goes up to more than 166 detections for the top 10% most observed sample that are part of the aforementioned high-cadence regions. On average, our SNe Ia are detected 12 d before maximum light and up to 55 d after with a median one day cadence in any filter. This corresponds to a typical 2.9, 2.5 and 6 d same-filter revisit for , and band, respectively.
Figure 1 illustrates typical worst, average and best sampling light curve of the release, with the increased uncertainty in the derived SALT2 values of and with decreasing sampling, shown in the legend.
@@
As detailed in Smith et al., the ZTF SN Ia DR2 light curves have been extracted using a custom re-calibration of the forced photometry pipeline presented in Masci et al. (2019) and Yao et al. (2019). The absolute zero-point, set to in the release light curves, is known at the 5 percent level, but the relative photometry is closer to 1%. Our limited knowledge of the photometry stems from the use of forced-photometry, since reference stars are not measurable in difference imaging. For the next ZTF SN Ia releases, we will use scene-modeling photometry (e.g., Holtzman et al., 2008; Astier et al., 2013; Brout et al., 2019) that do not require difference images to extract the transient light curve and thus enable to same flux estimator both on stars and on the transients of interest. We demonstrate in Lacroix et al. that we have a working scene-modeling pipeline, and by comparing it to the released DR2 light curves we can assert that our light curve colors (i.e., relative calibration) are good at the percent level; see details in Smith et al.. However, Lacroix et al. identified a non-linear bias in the flux measurements, on the per-cent level, dubbed the “pocket effect”. This effect was observed starting in November of 2019, when the camera wave-front read-out system characteristics had been updated. The origin of this effect, its modeling and its correction are presented in Lacroix et al., but the current ZTF light curves are affected by this read-out issue.
@@
We illustrate Fig. 2 (top) the typical impact of the pocket effect on our data: in comparison to their true magnitude, deviate by up to a few per-cent peak-to-peak. To test to impact to this effect on our results, we simulated realistic light curves, as in Amenouche et al. (see also Section 4.6), and we perturbed them with the non-linearity effect presented in the top panel of Fig. 2. We then derived their light curve parameters as presented in Section 4.7. The resulting shifts are shown in the bottom panels of Fig. 2. The pocket effect is expected to exacerbate fainter SNe Ia and bias their observed stretch standardization parameter. As expected, the light curve colors are not affected by the sensor issue. However, the stretch is recovered typically (70% of the average stretch error ) lower than what is given in input. This average shift is independent of the actually input value and could then be seen as a simple stretch zero-point definition bias. As discussed in Ginolin et al. (2024,a ZTFDR2) who study the stretch standardization, this stretch zero-point bias has no significant impact on our results since the absolute stretch value is meaningless as long as we do not compare the ZTF DR2 stretch values with these from another survey, as done, for instance, in Nicolas et al. (2021). In such a case, we suggest to apply a correction term. The SN Ia peak magnitude slowly drifts because of the pocket effect, especially beyond 18.5 mag, as shown in the panel of Fig. 2; so typically SN Ia at . But when focusing on the Hubble residuals, a similar trend than that of the stretch is observed: the zero point is off by (, when not accounting for the intrinsic scatter) and this shift is uncorrelated with the actual Hubble residuals value. Consequently, the pocket effect has no impact on pure-ZTF DR2 magnitude residual studies (e.g. steps), since this zero-point cancels out with cosmology ( definition). We emphasize that only targets acquired after Nov. 2019 are affected by the pocket effect. Hence, we encourage people using ZTF SN Ia DR2 data to test their results when only including the 50% of the ZTF SN Ia DR2 sample with peak magnitude prior the October 2019 as done, e.g., in Ginolin et al. (2024,a ZTFDR2).
@@
However, since the ZTF SN Ia DR2 light curves are derived from force-photometry and the above mentioned pocket effect, we discourage the use of this data release for precision cosmological parameter inference. The next release (DR2.5, planned for end-2025), will address both problems and hence will be suitable for precision cosmology, see Lacroix et al.. While it will remove a systematic bias, we nonetheless do not expect any significant reduction in the Hubble residuals dispersion once the pocket effect is corrected, given its negligible amplitude in comparison to the SN Ia intrinsic scatter.
@@ 4.2 Spectra
This DR2 contains 5138 spectra associated to the 3628 SNe Ia in our spectroscopically confirmed sample. Each target has at least one spectrum with sufficient quality to enable a secured “Ia” classification, and 28% have multiple spectra. Most of our spectra (60%) have been acquired by the SEDm and, thanks to the observing strategy of the BTS program, the vast majority of these have been acquired near or before maximum light. In addition to the SEDm, we obtained data from many other facilities, such as the Liverpool Telescope (7.6%; Steele et al., 2004), the Palomar 200-inch Hale Telescope and European Southern Observatory’s New Technology Telescope (NTT), as part of the ePESSTO program (6.5%; Smartt et al., 2015).
All spectra have been matched with the snid classification algorithm (Blondin & Tonry, 2007) using a custom template library made of 370 templates (available upon request). This matching was used as initial classification indicator and to derive SN-feature based redshifts for all our targets. For further details of how further sub-classifications were performed see Section 4.4, as well as Dimitriadis et al. and Burgaz et al. (2024,a ZTFDR2).
@@ 4.3 Host galaxies
As detailed in Smith et al., the host identification of each SN Ia in our sample is made in two steps. Firstly, we query public databases, such as DESI-LS DR9 (Dey et al., 2019), SDSS DR17 Abdurro’uf et al. (2022), and PS1 DR2 (Flewelling et al., 2020), for sources within an 100 kpc radius given the estimated redshifts from snid spectral template matching. Then, we compute the directional light radius (DLR, Sullivan et al., 2006; Gupta et al., 2016) between the SN and each surrounding galaxy or galaxy-like sources and identify the host as the source with the smaller DLR. SNe were no galaxy has a is excluded. The global photometry is derived using the HostPhot package (Müller-Bravo & Galbany, 2022) on public , , , , PS1 DR2 images that cover the same sky as ZTF. Local photometry is made using 2 kpc radius aperture photometry of these data (e.g. Briday et al., 2022).
@@
From the initial master list of 3668 targets having at least a ZTF light curve and a spectrum, we collectively visually inspected each target to ensure they were indeed a SN Ia and to sub-classify them when possible. This process has been made in two steps. Firstly, the ZTF Cosmology SWG members manually inspected the data through a specially designed web-application333typingapp.in2p3.fr. In the end, 32 users did more than 14,000 individual (sub)-classifications, that fed a decision-tree algorithm to automatically (sub)-classify the objects. This way, each SN Ia from the DR2 has been vetted at least twice, and 3.5 times on average.
In a second step, SN Ia population experts double-checked edge cases (including those with ambiguous sub-types suggested) in the volume-limited () sample to release the final (sub)-classification (see details in Dimitriadis et al. and Burgaz et al. 2024,a ZTFDR2). At the end of this procedure, about 5% of the SN Ia targets were classified as too peculiar to be included in cosmological analyses (e.g., “Ia-CSM” or “91bg-like”). However, some unidentified peculiar sub-types likely remain in the sample at that were not checked in detail. These contaminants are likely limited to overluminous events, such as “Ia-CSM” and “03fg-like” SNe Ia that would be detected at these higher redshift. These events are intrinsically rare and would also likely fail the SALT2 light curve quality cuts that are applied for cosmological measurements.
@@ 4.5 Redshift
Each of our 3628 targets has a redshift, that comes from one of four sources in order of preference: public galaxy redshift catalog (2200, 60.6%), galaxy emission lines visible in non-SEDm target spectra (199, 5.5%), galaxy emission lines visible in SEDm spectra (121, 3.3%), and estimated using snid SN Ia template matching (1086, 30.0%). The first two cases have a typical precision on the redshift, and are refereed to as “galaxy-redshifts”, while the latter have a precision of and are referred to has “SN-redshifts”. Of the 2200 galaxy catalog redshifts, 71% comes from the DESI MOST Hosts program (Soumagnac et al., 2024)
Figure 3 illustrates the accuracy and precision of the sn-feature based redshifts that account for 30% of the sample (22% for the volume limited sample ) using the sn-feature redshifts of SNe Ia with galaxy redshifts (mostly DESI, see Soumagnac et al., 2024). We conclude that sn-feature redshifts are unbiased (average shift lower than ) with a typical scatter of across the entire redshift range coverage by ZTF (see details in Smith et al.). Such a scatter () typically corresponds to a 0.09 mag additional scatter on the Hubble diagram at our median redshift .
Redshift are given in the heliocentric frame and the DR2 redshift distribution is shown in Fig. 5. Since most of our hosts are brighter than 20 mag in band, the vast majority of our SN-redshifts shall soon be acquired and released by DESI as part of their bright galaxy survey program (Hahn et al., 2023).
@@ 4.6 Selection
In Section 4 we claimed that our sample is volume limited up to . A fully realistic simulation based on observing logs analysis supporting this claim is presented in Amenouche et al.. In this section, we provide key elements characterizing the sample selection of our dataset.
As presented in Smith et al., the ZTF survey photometric depth is 20.4 mag in , 20.6 mag in and 20.0 mag in on average, but the spectroscopic follow-up is mag shallower. As we require a spectroscopic classification for the ZTF SN Ia DR2 release, this spectroscopic follow-up magnitude limit consequently is our limiting selection function. Figure 4 (top) presents our effective selection function, modeled as a survival sigmoid , with and . As expected, this model closely matches the BTS spectroscopic typing completeness measurements from Perley et al. (2020).
The steep fainter-redder SN Ia relation makes that the firsts SNe Ia missed by a magnitude limited survey are red targets (see e.g. discussion in Nicolas et al. 2021). Hence, the SN Ia color distribution is strongly sensitive to such selection, more so than, for instance, the SN Ia stretch. We show in Fig. 4 the color distribution of non-peculiar SNe Ia of the DR2 sample for various redshift ranges (, , , ). We over-plot our parent population model, finding it to be within , using a scatter derived from 3000 realistic simulations of the same number of targets than that observed in the DR2. For the parent population, we assume an intrinsic Gaussian distribution convoluted with an extrinsic exponential decay as given by Ginolin et al. (2024,b ZTFDR2) ; see also e.g., Brout & Scolnic (2021) and Popovic et al. (2024, ZTFDR2).
We draw two conclusions from this figure. First, since the simulations, which are sensitive to both the assumed parent population and the modeled selection function, closely match the observed data for all various redshift ranges, we conclude that our model for the selection and underlying SNIa distributions are reasonable. Then, second, since the observed SN Ia DR2 color distribution at closely matches the parent population (significant deviations are visible for the case), and since the color distribution is the most sensitive to selection effect, we conclude that our sample is not affected by significant selection function up to , i.e. it is volume-limited up to . See additional details in Amenouche et al..
@@ 4.7 Distances
We fit our light curves using the SALT2 (Guy et al., 2007, 2010) algorithm in its version 4 (Betoule et al., 2014) as retrained by Taylor et al. (2021) and made available by the sncosmo package444https://sncosmo.readthedocs.io/en/stable/about.html as “SALT2-T21”. As detailed in Smith et al., the fit is performed in the d rest-frame phase range as this is when the light curve algorithm is sufficiently trained (see also Rigault et al. 2024,b ZTFDR2). To do so the fit is performed twice. First, given a guess coming from the light curve data we cut at d and fit SALT2 to get a robust , then we apply the d cut to refit the light curve and store the SALT2 parameters, their errors and covariances.
@@
We show the SALT2 stretch () and color () parameters for all the 2667 SNe Ia passing our “basic quality” cuts in Fig 5. Above it can be seen that the fraction of red (high ) and slow (low ) SNe Ia starts to decrease rapidly due to selection effects since as these faster and redder SNe Ia are fainter. Below this redshift of 0.06, our sample is considered free from non-random selection functions for the normal SN Ia population (see section 4.6). We refer to this sample as “volume limited” and it contains nearly 1000 targets normal SNe Ia.
@@
We present our ZTF Hubble Diagram for the 2629 non-peculiar SNe Ia passing our quality cuts in Fig. 6. These SNe Ia have been standardized using the methodology detailed in Ginolin et al. (2024,a ZTFDR2), given the SALT2 stretch and color parameter and host local environmental properties (e.g., Sullivan et al., 2010; Rigault et al., 2020). Following that paper, we use , and (local-color step).
@@
Before standardization, our SNe Ia have a natural scatter along the Hubble diagram of 0.33 mag, using the normalized Median Absolute Deviation (nMAD) as robust scatter estimator. The standard deviation (std) is 0.46 mag (nMAD=std for a normal distribution). This scatter reduces down to 0.165 mag after standardization (std: 0.209 mag) accounting for all 2629 ZTF DR2 non-peculiar SNe Ia.
@@
A part of this scatter is due to peculiar motions that typically are km s-1, leading to an additional scatter of 0.03 mag at our median redshift (), but to more than 0.1 mag for targets with as illustrated in Fig. 6. Another part is caused by the use of sn-feature redshifts for 32% of the sample. As presented in Section 4.5, these have a dispersion of 900 km s-1 corresponding to an additional Hubble residual scatter of mag at . To estimate our Hubble scatter without the impact of peculiar velocities and sn-feature redshifts, we take the scatter in the range of . Their the standardized SN Ia scatter is 0.150 mag (nMAD) corresponding to 0.145 mag after removing the expected peculiar velocity contribution. This scatter is illustrated in the inset panel of Fig. 6, showing it is a good description of our 2629 SN Ia standardize Hubble residual dispersion. Such 0.15 magnitude scatter is in good agreement with the state-of-the-art results from the literature (see e.g., Brout et al., 2022).
@@
We finally inspect for outliers that may affect our standard deviation (std) measurements (nMAD statistics been robust). Given our sample size, the Chevaunet’s criteria (less that 50% of chance to detect such an outlier) corresponds to a rejection. Doing so removes 23 objects (0.8% of the total sample) and leads to a total standard deviation (std) of (nMAD: 0.159 mag), including 2606 SN Ia, and (nMAD: 0.144 mag) when discarding sn-feature redshift SNe Ia (1647 left, rejection).

### v2

@@ 4 Overview of the released data
Details concerning the DR2 data and the derived parameters are given in Smith et al. (in prep.). This section summarizes the top-level information.
@@ 4.1 Light curves
In this release, we provide , , and ZTF light curves for all our 3628 SNe Ia acquired by the ZTF camera installed on the P48. Additional photometric data points, such as those obtained from the SED machine in camera mode, are not included in this release.
@@
On the 2960 SNe Ia passing the good sampling criteria, only 9 do not have -band and only 2 do not have -band detections in the rest-frame phase range of 10 to +40 with respect to maximum light. However, only 46% of our targets have -band detections in this same phase range. Most targets acquired since early 2021 will have coverage in all three bands as we updated our observing strategy, but they are not part of this sample. Future releases of DR2 targets will include additional -band observations, after the remaining reference images are processed. For these 2960 targets, we typically have 40 detections (median) in the day phase range, with medians of 9 and 27 points pre- and post-maximum light, respectively. This goes up to more than 166 detections for the top 10% most observed sample that is part of the aforementioned high-cadence regions. On average, our SNe Ia are detected 12 d before maximum light and up to 55 d after, with a median one-day cadence in any filter. This corresponds to a typical 2.9, 2.5, and 6 d same-filter revisit for the , , and bands, respectively. Figure 1 illustrates the typical worst, average, and best sampling light curve of the release, with the increased uncertainty in the derived SALT2 values of and with decreasing sampling, shown in the legend.
@@ 4.1.2 Photometry
As is detailed in Smith et al. (in prep.), the ZTF SN Ia DR2 light curves have been extracted using a custom recalibration of the forced-photometry pipeline presented in Masci et al. (2019) and Yao et al. (2019). The absolute zero-point, set to in the released light curves, is known at the 5 percent level, but the relative photometry is closer to 1%. Our limited knowledge of the photometry stems from the use of forced photometry, since reference stars are not measurable in difference imaging. For the next ZTF SN Ia releases, we shall use scene-modeling photometry (e.g., Holtzman et al., 2008; Astier et al., 2013; Brout et al., 2019) that do not require difference images to extract the transient light curve, and thus enable one to use the same flux estimator on both stars and the transients of interest. We demonstrate in Lacroix et al. (in prep.) that we have a working scene-modeling pipeline, and by comparing it to the released DR2 light curves we can assert that our light curve colors (i.e., relative calibration) are good at the percent level (see details in Smith et al. in prep.). However, Lacroix et al. (in prep.) identified a nonlinear bias in the flux measurements, on the percent level, dubbed the “pocket effect.” This effect was observed starting in November of 2019, when the camera wave-front read-out system characteristics had been updated. The origin of this effect, its modeling, and its correction are presented in Lacroix et al. (in prep.), but the current ZTF light curves are affected by this read-out issue.
@@
We illustrate in Fig. 2 (top) the typical impact of the pocket effect on our data: in comparison to their true magnitudes, they deviate by up to a few percent peak to peak. To test the impact of this effect on our results, we simulated realistic light curves, as in (Amenouche et al., 2024, see also our Sect. 4.6), and we perturbed them with the nonlinearity effect presented in the top panel of Fig. 2. We then derived their light curve parameters, as is presented in Sect. 4.7. The resulting shifts are shown in the bottom panels of Fig. 2. The pocket effect is expected to exacerbate fainter SNe Ia and bias their observed stretch standardization parameter. As was expected, the light curve colors are not affected by the sensor issue. However, the stretch recovered is typically (70% of the average stretch error, ) lower than what is given as an input. This average shift is independent of the actual input value and could be seen as a simple stretch zero-point definition bias. As is discussed by Ginolin et al. (2024,a), who study the stretch standardization, this stretch zero-point bias has no significant impact on our results, since the absolute stretch value is meaningless as long as we do not compare the ZTF DR2 stretch values with these from another survey, as was done, for instance, in Nicolas et al. (2021). In such a case, we suggest applying a correction term.
The SN Ia peak magnitude slowly drifts because of the pocket effect, especially beyond 18.5 mag, as is shown in the panel of Fig. 2; so typically SN Ia at . But when focusing on the Hubble residuals, a similar trend to that of the stretch is observed: the zero point is off by (, when not accounting for the intrinsic scatter) and this shift is uncorrelated with the actual Hubble residuals value. Consequently, the pocket effect has no impact on pure ZTF DR2 magnitude residual studies (e.g., steps), since this zero-point cancels out with cosmology ( definition).
@@
We emphasize that only targets acquired after November 2019 are affected by the pocket effect. Hence, we encourage people using ZTF SN Ia DR2 data to test their results when only including the 50% of the ZTF SN Ia DR2 sample with a peak magnitude prior to October 2019 as was done, for example, in Ginolin et al. (2024,a). However, since the ZTF SN Ia DR2 light curves are derived from forced photometry and the abovementioned pocket effect, we discourage the use of this data release for precision cosmological parameter inference. The next release (DR2.5, planned for the end of 2025), will address both problems, and hence will be suitable for precision cosmology (see Lacroix et al. in prep.). While it will remove a systematic bias, we nonetheless do not expect any significant reduction in the Hubble residuals dispersion once the pocket effect is corrected, given its negligible amplitude in comparison to the SN Ia intrinsic scatter.
@@ 4.2 Spectra
This DR2 contains 5138 spectra associated with the 3628 SNe Ia in our spectroscopically confirmed sample. Each target has at least one spectrum with sufficient quality to enable a secured “Ia” classification, and 28% have multiple spectra. Most of our spectra (60%) have been acquired by the SEDm and, thanks to the observing strategy of the BTS program, the vast majority of these have been acquired near or before maximum light. In addition to the SEDm, we obtained data from many other facilities, such as the Liverpool Telescope (7.6%; Steele et al., 2004), the Palomar 200-inch Hale Telescope , and European Southern Observatory’s New Technology Telescope (NTT), as part of the ePESSTO program (6.5%; Smartt et al., 2015).
All of the spectra were matched with the snid classification algorithm (Blondin & Tonry, 2007) using a custom template library made of 370 templates (available upon request). This matching was used as an initial classification indicator and to derive SN-feature-based redshifts for all our targets (for further details of how further subclassifications were performed see Sect. 4.4, as well as Dimitriadis et al. (2024) and Burgaz et al. (2024,b)).
@@ 4.3 Host galaxies
As is detailed in Smith et al. (in prep.), the host identification of each SN Ia in our sample was made in two steps. First, we queried public databases, such as DESI-LS DR9 (Dey et al., 2019), SDSS DR17 Abdurro’uf et al. (2022), and PS1 DR2 (Flewelling et al., 2020), for sources within an 100 kpc radius, given the estimated redshifts from snid spectral template matching. Then, we computed the directional light radius (DLR, Sullivan et al., 2006; Gupta et al., 2016) between the SN and each surrounding galaxy or galaxy-like source and identified the host as the source with the smaller DLR. The SNe for which no galaxy has a were excluded. The global photometry was derived using the HostPhot package (Müller-Bravo & Galbany, 2022) on public , , , , PS1 DR2 images that cover the same sky as ZTF. Local photometry was made using the 2 kpc radius aperture photometry of these data (e.g. Briday et al., 2022). Once optical global and local photometry had been estimated, we computed stellar masses and rest-frame color using a spectral energy distribution fitting performed with the PÈGASE2 galaxy spectral templates (Le Borgne & Rocca-Volmerange, 2002), as in, for example, Sullivan et al. (2010).
@@
From the initial master list of 3668 targets that have at least a ZTF light curve and a spectrum, we collectively visually inspected each target to ensure they were indeed an SN Ia and to subclassify them when possible. This process was made in two steps. Firstly, the ZTF Cosmology SWG members manually inspected the data through a specially designed web application.333typingapp.in2p3.fr In the end, 32 users did more than 14,000 individual (sub)-classifications, which fed a decision-tree algorithm to automatically (sub)-classify the objects. This way, each SN Ia from the DR2 was vetted at least twice, and 3.5 times on average.
In a second step, SN Ia population experts double-checked edge cases (including those for which ambiguous subtypes were suggested) in the volume-limited () sample to release the final (sub)-classification (see details in Dimitriadis et al. 2024 and Burgaz et al. 2024,b). At the end of this procedure, about 5% of the SN Ia targets were classified as too peculiar to be included in cosmological analyses (e.g., “Ia-CSM” or “91bg-like”). However, some unidentified peculiar subtypes likely remain in the sample at that were not checked in detail. These contaminants are likely limited to overluminous events, such as “Ia-CSM” and “03fg-like” SNe Ia that would be detected at these higher redshifts. These events are intrinsically rare and would also likely fail the SALT2 light curve quality cuts that are applied for cosmological measurements.
@@ 4.5 Redshift
Each of our 3628 targets has a redshift that comes from one of four sources in order of preference: the public galaxy redshift catalog (2200, 60.6%), galaxy emission lines visible in non-SEDm target spectra (199, 5.5%), galaxy emission lines visible in SEDm spectra (121, 3.3%), and estimated using snid SN Ia template matching (1086, 30.0%). The first two cases have a typical precision on the redshift, , and are referred to as “galaxy-redshifts,” while the latter have a precision of and are referred to as “SN redshifts.” Of the 2200 galaxy catalog redshifts, 71% come from the DESI MOST Hosts program (Soumagnac et al., 2024).
Figure 3 illustrates the accuracy and precision of the SN-feature based redshifts that account for 30% of the sample (22% for the volume-limited sample, ) using the SN-feature redshifts of SNe Ia with galaxy redshifts (mostly DESI, see Soumagnac et al., 2024). We conclude that SN-feature redshifts are unbiased (average shifts lower than ) with a typical scatter of across the entire redshift range coverage by ZTF (see details in Smith et al. in prep). Such a scatter () typically corresponds to a 0.09 mag additional scatter on the Hubble diagram at our median redshift of .
Redshift are given in the heliocentric frame and the DR2 redshift distribution is shown in Fig. 5. Since most of our hosts are brighter than 20 mag in the band, the vast majority of our SN redshifts will soon be acquired and released by DESI as part of their bright galaxy survey program (Hahn et al., 2023).
@@ 4.6 Selection
In Sect. 4, we claimed that our sample is volume-limited up to . A fully realistic simulation based on observing log analysis supporting this claim is presented in Amenouche et al. (2024). In this section, we provide key elements characterizing the sample selection of our dataset.
As is presented in Smith et al. in prep., the ZTF survey photometric depth is 20.4 mag in , 20.6 mag in and 20.0 mag in on average, but the spectroscopic follow-up is mag shallower. As we require a spectroscopic classification for the ZTF SN Ia DR2 release, this spectroscopic follow-up magnitude limit consequently is our limiting selection function. Figure 4 (top) presents our effective selection function, modeled as a survival sigmoid, , with and . As was expected, this model closely matches the BTS spectroscopic typing completeness measurements from Perley et al. (2020).
The steep fainter-redder SN Ia relation means that the first SNe Ia missed by a magnitude-limited survey are red targets (see e.g., discussion in Nicolas et al. 2021). Hence, the SN Ia color distribution is highly sensitive to such a selection, more so than, for instance, the SN Ia stretch. We show in Fig. 4 the color distribution of non-peculiar SNe Ia of the DR2 sample for various redshift ranges (, , , and ). We overplot our parent population model, finding it to be within , using a scatter derived from 3000 realistic simulations of the same number of targets than that observed in the DR2. For the parent population, we assume an intrinsic Gaussian distribution convolved with an extrinsic exponential decay given by Ginolin et al. (2024,b) (see also e.g., Brout & Scolnic (2021) and Popovic et al. (2024)).
We draw two conclusions from this figure. First, since the simulations, which are sensitive to both the assumed parent population and the modeled selection function, closely match the observed data for all of the various redshift ranges, we conclude that our model for the selection and underlying SNIa distributions are reasonable. Second, since the observed SN Ia DR2 color distribution at closely matches the parent population (significant deviations are visible in the case), and since the color distribution is the most sensitive to selection effects, we conclude that our sample is not affected by a significant selection function up to ; that is, it is volume-limited up to (see additional details in Amenouche et al. (2024)).
@@ 4.7 Distances
We fit our light curves using the SALT2 (Guy et al., 2007, 2010) algorithm in its version 4 (Betoule et al., 2014) retrained by Taylor et al. (2021) and made available by the sncosmo package444https://sncosmo.readthedocs.io/en/stable/about.html as “SALT2-T21.” As is detailed in Smith et al. in prep, the fit was performed in the d rest-frame phase range as this is when the light curve algorithm is sufficiently trained (see also Rigault et al. 2024,b). To do so, the fit was performed twice. First, given a guess coming from the light curve data, we cut at d and fit SALT2 to get a robust , then we applied the d cut to refit the light curve and store the SALT2 parameters, their errors, and covariances.
@@
We show the SALT2 stretch () and color () parameters for all the 2667 SNe Ia passing our “basic quality” cuts in Fig 5. Above , it can be seen that the fraction of red (high ) and slow (low ) SNe Ia starts to decrease rapidly due to selection effects, since these faster and redder SNe Ia are fainter. Below this redshift of 0.06, our sample is considered to be free from nonrandom selection functions for the normal SN Ia population (see Sect. 4.6). We refer to this sample as “volume limited” and it contains nearly 1000 normal SN Ia targets.
@@
We present our ZTF Hubble diagram for the 2629 non-peculiar SNe Ia that pass our quality cuts in Fig. 6. These SNe Ia have been standardized using the methodology detailed in Ginolin et al. (2024,a), given the SALT2 stretch and color parameter and host local environmental properties (e.g., Sullivan et al., 2010; Rigault et al., 2020). Following that paper, we have used , , and (local-color step).
@@
Before standardization, our SNe Ia have a natural scatter along the Hubble diagram of 0.33 mag, using the normalized median absolute deviation (nMAD) as a robust scatter estimator. The standard deviation (std) is 0.46 mag (nMAD=std for a normal distribution). This scatter reduces down to 0.165 mag after standardization (std: 0.209 mag), accounting for all 2629 of the ZTF DR2 non-peculiar SNe Ia.
@@
Part of this scatter is due to peculiar motions that typically are km s-1, leading to an additional scatter of 0.03 mag at our median redshift (), but to more than 0.1 mag for targets with , as is illustrated in Fig. 6. Another part is caused by the use of SN-feature redshifts for 32% of the sample. As is presented in Sect. 4.5, these have a dispersion of 900 km s-1, corresponding to an additional Hubble residual scatter of mag at . To estimate our Hubble scatter without the impact of peculiar velocities and SN-feature redshifts, we took the scatter in the range of . There, the standardized SN Ia scatter is 0.150 mag (nMAD), corresponding to 0.145 mag after removing the expected peculiar velocity contribution. This scatter is illustrated in the inset panel of Fig. 6, showing it is a good description of our 2629 SN Ia standardized Hubble residual dispersion. Such a 0.15 magnitude scatter is in good agreement with the state-of-the-art results from the literature (see e.g., Brout et al., 2022).
@@
We finally inspected for outliers that may affect our standard deviation (std) measurements (nMAD statistics have been robust). Given our sample size, Chevaunet’s criteria (a less than 50% chance of detecting such an outlier) correspond to a rejection. Doing so removed 23 objects (0.8% of the total sample) and led to a total standard deviation (std) of (nMAD: 0.159 mag), including 2606 SN Ia, and (nMAD: 0.144 mag) when discarding SN-feature redshift SNe Ia (1647 left, rejection).